OpenStudy (anonymous):

Find the standard form of the equation of the parabola with a focus at (0, 4) and a directrix at y = -4.

OpenStudy (anonymous):

hey :)

OpenStudy (skullpatrol):

hi :)

OpenStudy (anonymous):

do you know how to do this problem?

OpenStudy (skullpatrol):

Isn't there a standard formula?

OpenStudy (anonymous):

(x - h)^2 = 4p(y - k)

OpenStudy (anonymous):

i tried doing it but i got y=1/32 * (x-4)^2 - 3

OpenStudy (anonymous):

some weird answer

OpenStudy (skullpatrol):

Now just multiply it out.

OpenStudy (anonymous):

did you get that too?

OpenStudy (skullpatrol):

what do h, k and p stand for?

OpenStudy (anonymous):

h and k is the vertex

OpenStudy (anonymous):

p= k-y so p =8

OpenStudy (jdoe0001):